Three moles of helium gas at 300°C and 2.0 bar are expanded isothermally against a constant external pressure of 0.25 bar. Determine the quantities q , w , ΔU , ΔSsys , and ΔGsys for this process.
1
Expert's answer
2020-10-08T13:54:52-0400

Three moles of helium gas at 300°C and 2.0 bar are expanded isothermally .


So , Volume of gas ( initially ) = 3 * 0.0821 * 573 / 2 = 70.5 liter .


Volume of gas ( final ) . = 3 * 0.0821 * 573 / 0.25 = 564.52 liter .


SO , WORK DONE = W = - 0.25 ( 564.52 - 70.5 ) * 101.2 j = - 12498.706 j . = - 12.498 kj .


SO , q = - W = + 12.498 kj


SO ,  ΔU = ZERO .


SO , ΔSsys = q / T = 12.498 * 1000 / 573 = 21.811 j / k .mol


ΔHsys = ZERO .


SO , ΔGsys = ΔHsys - T . ΔSsys = 0 - 573 * 21.811 = 12497.8 2 == 12.5 KJ / MOL .